OpenStudy (anonymous):

y=1/(x^2-2)^3 diferentiation help plzzz

OpenStudy (anonymous):

This would be chain rule do you know it?

OpenStudy (anonymous):

expound on it plzz

OpenStudy (anonymous):

Chain rule if i have a function inside a function f(g(x)) The derivative of such a function would be (f'(g(x)) g'(x) Example if you want to dufferentiaite log( x^2) Here f(x)=logx and g(x)=x^2. So first find f'(x)=1/x so f'(g(x))=1/x^2 g'(x)=2x. So the net derivative is f'(g(x)) g'(x)= 2x/x^2=2/x.

OpenStudy (anonymous):

Did you understand till here?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

Ok so here f(x)=1/x^3 and g(x)=x^2-2. Try to do it by yourself... Basically find f' and then replace x by g(x). Multiply whatever you got by g'

OpenStudy (anonymous):

Did you get it?